Solid silver acetate is slowly added to 75.0 mL of a 0.0383 M sodium bromide solution. The concentration of silver ion required to just initiate precipitation is _______ M. 
Answer #1

Ag+(aq) + Br-(aq) ----> AgBr(s)

Ksp of AgBr = [Ag+][Br-]

Ksp of AgBr = 4.1*10^-13

[Ag+] = ?

[Br-] = 0.0383 M

4.1*10^-13 = x* 0.0383

x = 1.07*10^-11

concentration of Ag+ required to initiate preipitate = 1.07*10^-11 M